# Distributed Multisensor ISAC

### v1

@@ V-E Sensing Cycles – OFDMA Resource Allocation, Time-frequency Precoding and Link Adaptation
This finally leads us to the question of whether a similar focusing technique is possible in Doppler domain. Bringing the observed Doppler shift to zero for a moving target would support a longer coherent integration time for correlation processing. This can be important for target detection and reliable tracking and again follows the idea of shifting the processing effort from the sensor as “precoding” to the infrastructure. The result can be seen as a 2D channel matched filter that focuses the received signal in the joint 2D delay-Doppler domain. Obviously, focusing on the time delay as described above will also impact the Doppler domain, since the delay corresponds to range and the Doppler shift corresponds to speed, i.e., the range derivative. Therefore, if mirroring in delay works perfectly for range focusing, the Doppler contribution by the different paths would also collapse to zero. However, in practical situation this might not be so easy. This is because the delay resolution and variance is limited by the bandwidth and noise constraint of each OFDM symbol. Doppler, however, relates to carrier phase, which is much more sensitive to range variance. In 2D delay-Doppler estimation, Doppler by Fourier transforming along the slow time Figure 15, which may include several hundred or even more consecutive OFDM symbols. Phase stability along the coherent processing interval is of highest importance to get a stable Doppler shift that describes a linear movement. While FFT filtering along the slow time reduces the variance, the immanent variance of the channel estimates used for the TR operation can actually worsen the situation. Therefore, model based joint delay-Doppler estimation should work better as the inherent model would cause some smoothing along slow time. Moreover, since the model order can be controlled, this would pave the way towards using only the dominating and persisting paths for delay-Doppler precoding, e.g. those paths that are stable and reflected by objects at know positions, which can be traced back for explicit multipath exploiting localization.
@@
Regarding the influence on the target object we can deduce from Figure 21 that the diversity advantage for increased target detection probability applies here as well, as was discussed for distributed MS ISAC. However, the geometric advantage for localization only applies if we can use the true positions of scatterers for explicit, model-based localization. Here, the circle closes and it becomes clear that TRF and ray optical modelling of propagation can be combined. However, recalling the special role of the target as a solitary interacting object in the propagation environment (see Section 4, ’Multipath Propagation’), it deserves even more attention for waveform adaptation. While delay-Doppler precoding can be understood as a 2D channel-matched filter, a similar and equivalent matching operation can be assigned to the target. Without going into every detail, we will summarize some aspects. For extended targets (where the radar bandwidth allows resolving several delay bins that are related to the same target) it may make sense to adapt the illuminating signal to the target response by time reversal. The result can be considered a target-matched filter that maximizes coherently reflected energy while avoiding its spread to several delay bins. This seems to be a better metric (because it is ultimate) for calculating the radar link budget than the radar cross-section, which presumes non-extended targets.
@@
It seems that a wider bandwidth can resolve more structural details about the target, which would also favor higher frequencies. But this could be a serious pitfall since besides some material-related frequency dependency of reflectivity, there may also be a strong structural dependency of reflectivity. This is very often related to cavities being a part of the target structure. Then the target can respond with strong Eigenmodes if properly exited. Since Eigenmodes are related to resonance effects, the proper frequency of the illumination signal is crucial. So we have to make sure that the resonance frequency is covered. Depending of the size of interesting targets, these Eigenfrequencies can be quite low (e.g. FR1). Therefore, multiband target illumination as described in the subsection on Heterogonous Radio Nodes can be useful.
@@
Another domain for target response estimation and adaptation is polarization. According to the Jones matrix calculus we can estimate the $2\times 2$ polarimetric target response matrix by illumination with two orthogonal polarizations (dual linear or circular) and sensing of the scattered waves in orthogonal polarization orientations [ref68]. The input/output Jones vector is defined by the dual polarimetric antennas at Tx and Rx. In general, orthogonal access at the two Tx antenna ports (e.g. by TDMA) would be appropriate. Depending on its shape, the reflectivity of the target may strongly depend on the polarization orientation of the illuminating waves and somehow the target structure acts as a more or less polarization sensitive filter. This may require adaptation of polarization to maximize target reflectivity. The polarization orientation of the illuminating wave can be controlled by proper weighting and simultaneous transmission of the waveform from two polarimetric Tx antenna ports. With phase control included, also circular or elliptic polarization can be generated from two orthogonal linear polarized antenna port. The same adaptation principle also applies to the Rx chain. Note that usual definitions of linear polarization as Vertical (V) and Horizontal (H) applies to the local coordinate system of Tx and Rx antennas, which may be not aligned in H and V between each other.The same applies to the local coordinate system of the target and also does not apply to the Earth-centered coordinate system. This means that just simplifying to vertical (VV) or horizontal (HH) polarization at the transmitter (Tx) and receiver (Rx) is not enough in most cases if the Tx and Rx antennas are randomly oriented in space as it may be the case if agile ISAC sensor platforms are included.
@@ V-F Spatial Precoding, Beamforming, and Bidirectional Estimation
Spatial precoding is a well established method in mobile radio that came into play with the introduction of multiple antennas. It helps increase link performance, reduce interference, host and address multiple users, and increase data rate and bandwidth efficiency. Spatial precoding will also be very important for ISAC, especially for MS ISAC. Since 5G NR uses a flexible architecture based on codebook and non-codebook (full adaptive) schemes, spatial radio access can be controlled for omnidirectional or directive sensing. Without going too much to the technical details we should emphasize that spatial precoding, especially beamforming may be more disposable at higher frequencies (FR2) where hybrid (mixed analog and digital) antenna interfaces are applied. Multiple simultaneous beams for the same user are also possible. 5G NR allows transmit precoding and receive combining and also supports beamforming in the UL, which means that the gNB can assign a spatial precoding matrix to the UE, which can be a dedicated device acting as a sniffer sensor, as described above. In the sequel, we will discuss some striking applications of beamforming and spatial access for MS ISAC, though we do not claim to be exhaustive.
@@
The initial purpose of beamforming is to enhance the SNR (and consequently the range), as it is associated with an antenna gain. In the context of distributed cooperative MS ISAC networks, one application is to balance the SNR to achieve smooth coverage and performance distribution irrespective of the spatial density of the sensors. MS ISAC network densification will also go along with a proper choice of the frequency. Higher frequencies will be more appropriate at sensing hotspots since the wider bandwidth at FR2 offers more resolution capabilities in delay. But, because of the Friies equation, higher frequencies are more subjected for transmission loss since the effective antenna area is becoming smaller for the same gain. This can be well compensated by antenna arrays. However, this is accompanied by directivity. Hence, we would need time-consuming angular scanning in the surveillance cycle. Multifrequency sensing can help to get a fast (quick and dirty) overview picture of the scene and some first indication of possible targets using omnidirectional antennas a lower frequencies that does not require expensive angular scanning. The result is then used to instruct the FR2 beamformer in a second step [ref83].
@@
Beamforming directivity can be applied deliberately to reduce interference and multipath clutter through spatial filtering, especially when executed in a bidirectional manner at both the transmitter (Tx) and receiver (Rx). The antenna architecture depicted in Figure 22 is circular on both sides, making it ideal for symmetric D2D scenarios, such as V2V, with arbitrary spatial orientation of the devices. An equivalent antenna setup can be found for I2V communication, where planar arrays at the infrastructure side maybe more appropriate. The direct line-of-sight (LOS) link is important for data transfer and for obtaining the correlation reference. LOS beamforming suppresses multipath, which supports transmit signal recovery. A second surveillance channel points to the target, while clutter 1 is not detected by both beams and is therefore not visible. The case of target/clutter 2 is particularly interesting. Reflections from this object could not be filtered out by the Tx beam, but can be well suppressed by the Rx beam pointing towards Tx. Therefore, the Rx-LOS beam is important for receiving a clean, multipath-free copy of the transmitted signal. When the same object is considered as the target, the corresponding Rx beam pointing to the target makes it visible and also reduces LOS interference. This example shows, how the two bidirectional beams support each other and may increase directive filtering a lot. It is worth mentioning that, in a 3D geometry, matching the two opposite beams must also include polarization orientation, especially when azimuth and elevation matter and the orientation of the antennas may be more or less arbitrary [ref68]. Since the two beams are served by the same radio interface, mutual coherence is maintained, which is important for eTof and eDoppler estimation in CPCL.
@@
Figure 22: Bidirectional beamforming supporting clean LOS filtering and clutter supression
@@
Although the 5G beamforming interface hides the antenna in/outputs for individual access and processing, high resolution DoD/DoA estimation can still be achieved by smart use of the precoding vectors. In case of hybrid beamforming, the outputs of the analog multibeam network maps a huge number of hundreds or even thousands physical antennas to a small number of virtual antenna ports where digitization takes place. [ref97] describes beamformer design for DoD estimation based on limited-resolution phase shifters. However, as discussed above regarding high-resolution angular estimation, there are other approaches that allow for true high-resolution direction finding which goes beyond the mere resolution of beamforming. We will shortly introduce two approaches. Both solutions have in common that more than one beam is generated, that the beam patterns overlap, and that the multiple beam outputs explicitly differ in phase. The required number of virtual antenna ports is relatively small. This is because we only need to resolve those paths by angle that are not resolved by other parameters (delay, Doppler, or the complementary direction).
@@
A simple example is given in Figure 23. There are two beams formed, both of which have the same gain pattern. The beamforming vector is effectively shifted by one antenna element, which emulates two identical, overlapping subarrays effectively shifted by one antenna element. This results in translation-invariant radiation patterns meaning that any impinging wave is weighted by the same gain pattern. However, as the two beams have different phase centers, they are very well suited for high-resolution DoA estimation. The phase difference is the decisive factor for high-resolution DoA estimation, much more so than magnitude as it is the case with just beamforming. The well-known monopulse technique for high-resolution DoA estimation in radar is reminiscent of a simple replica of the same principle in the analog world. On a more recent time scale, the two arrays directly correspond to the shift-invariant pair of arrays that forms the basic principle of the ESPRIT algorithm. Obviously, the proposed approach can be further modified by increasing the number of subarrays and varying the phase center distances. Multiple overlapping shift-invariant subarrays would allow resolving multiple wave directions by subarray smoothing. This method can be considered as a hybrid implementation of the Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT) algorithm. It is limited to linear and planar arrays, which is not a serious draw back since integrated arrays primarily come as planar structures. Another advantage is that is does not require a specific array design with very sharp pencil beams. Instead, DoA estimation accuracy is gained by phase comparison. Therefore it is more important that the phase shifters are identical. Resolution of multiple paths within the same common beam is possible by exploiting the phase differences. On the other hand, beamforming can suppress waves from outside of the uniquely resolvable sector. This is indicated in Figure 23 by the more narrow beam. This would allow the distance between antenna elements to be extended beyond the half-lambda limit, which would increase estimation accuracy.
@@
Even if the goal is to estimate the angle based on maximum likelihood, the radiation pattern design can be relaxed since the specific pattern is not as important. However, we need to know how the complex pattern transforms from the specified coefficients to the radiated field in order to enable the iterative cancellation of overlapping beam patterns. Thus, the focus shifts from optimizing quantized coefficients to calibrating the radiation pattern. Lastly, estimation of DoD requires identifying the waveforms transmitted via the different antenna ports at the receive (Rx) side. Therefore, some kind of orthogonalization in the frequency-time frame is necessary.
@@
Figure 23: Emulating two shift-invariant radiation patterns for high resolution direction finding. The wider beams allows to cover a wide angular sector, while the more narrow beams reduces the angular field of view by spatial filtering.
@@
The multidimensionality aspect of parameter estimation refers to the bistatic propagation scenario, which includes the bidirectional geometry (in azimuth and elevation at Tx and Rx), ToF and Doppler for each propagation path, weighted with a $2\times 2$ complex-valued polarimetric Jones matrix. Thus, the bidirectional estimation extends the ToF-Doppler estimation described in the subsection on Model-based Bistatic Target Parameter Estimation to a joint 6D estimation procedure. When performing DoA/DoD estimations, impairments of antennas and antenna arrays can be of highest importance as they are most susceptible to RF effects. In this case, the device data model will be determined by calibration measurements of the installed radiation patterns. The traditional approach of model based estimation decomposes in a local gradient based and a global parameter optimization procedure with constraints. Actual research challenges address efficient real-time implementation in networks with distributed computational resources and AI support to multi-domain convex optimization with detection of relevant propagation paths. A related issue is in situ training, which effectively updates calibration throughout operation [ref67], [schieler2025prop_est_cnn]. Finally, advanced AI-based methods can be used to adjust the model parameters, as this can be a challenging problem of non-convex optimization, which can become even more complex when device data from calibration are not perfect or not fully available from calibration measurements in reference setups [ref35], but have to be learned implicitly by ML in operational situations.

### v2

@@ V-E Sensing Cycles – OFDMA Resource Allocation, Time-frequency Precoding and Link Adaptation
Doppler precoding. This finally leads us to the question of whether a similar focusing technique is possible in the Doppler domain. Bringing the observed Doppler shift to zero for a moving target would support a longer coherent integration time for correlation processing. This can be important for target detection and reliable tracking, and again follows the idea of shifting the processing effort from the sensor as “precoding” to the infrastructure. The result can be seen as a 2D channel matched filter that focuses the received signal in the joint 2D delay-Doppler domain. Obviously, focusing on the time delay as described above will also impact the Doppler domain, since the delay corresponds to range and the Doppler shift corresponds to speed, i.e., the range derivative. Therefore, if mirroring in delay works perfectly for range focusing, the Doppler contribution by the different paths would also collapse to zero. However, in a practical situation, this might not be so easy. This is because the delay resolution and variance are limited by the bandwidth and noise constraint of each OFDM symbol. Doppler, however, relates to the carrier phase, which is much more sensitive to range variance. In 2D delay-Doppler estimation, Doppler is obtained by FT along the slow time Figure 15, which may include several hundred or even more consecutive OFDM symbols. Phase stability along the coherent processing interval is of highest importance to get a stable Doppler shift that describes a linear movement. While FFT filtering along the slow time reduces the variance, the inherent variance of the channel estimates used for the TR operation can actually worsen the situation. Therefore, model-based joint delay-Doppler estimation should work better as the inherent model would cause some smoothing along slow time. Moreover, since the model order can be controlled, this would pave the way towards using only the dominating and persisting paths for delay-Doppler precoding, e.g., those paths that are stable and reflected by objects at known positions, which can be traced back for explicit multipath exploiting localization.
@@
Target related precoding. Regarding the influence on the target object, we can deduce from Figure 21 that the diversity advantage for increased target detection probability applies here as well, as was discussed for distributed MS-ISAC. However, the geometric advantage for localization only applies if we can use the true positions of scatterers for explicit, model-based localization. Here, the circle closes, and it becomes clear that TRF and ray optical modeling of propagation can be combined. However, recalling the special role of the target as a solitary interacting object in the propagation environment (see Section 4, ’Multipath Propagation’), it deserves even more attention for waveform adaptation. While delay-Doppler precoding can be understood as a 2D channel-matched filter, a similar and equivalent matching operation can be assigned to the target. Without going into every detail, we will summarize some aspects. For extended targets (where the radar bandwidth allows resolving several delay bins that are related to the same target), it may make sense to adapt the illuminating signal to the target response by time reversal. The result can be considered a target-matched filter that maximizes coherently reflected energy while avoiding its spread to several delay bins. This seems to be a better metric (because it is ultimate) for calculating the radar link budget than the radar cross-section, which presumes non-extended targets.
@@
It seems that a wider bandwidth can resolve more structural details about the target, which would also favor higher frequencies. But this could be a serious pitfall since, besides some material-related frequency dependency of reflectivity, there may also be a strong structural dependency of reflectivity. This is very often related to cavities being a part of the target structure. Then the target can respond with strong Eigenmodes if properly excited. Since Eigenmodes are related to resonance effects, the proper frequency of the illumination signal is crucial. So we have to make sure that the resonance frequency is covered. Depending on the size of interesting targets, these Eigenfrequencies can be quite low (e.g., FR1). Therefore, multiband target illumination as described in the subsection on Heterogeneous Radio Nodes can be useful.
@@
Another domain for target response estimation and adaptation is polarization. According to the Jones matrix calculus, we can estimate the $2\times 2$ polarimetric target response matrix by illumination with two orthogonal polarizations (dual linear or circular) and sensing of the scattered waves in orthogonal polarization orientations [ref68]. The input/output Jones vector is defined by the dual polarimetric antennas at Tx and Rx. In general, orthogonal access at the two Tx antenna ports (e.g., by TDMA) would be appropriate. Depending on its shape, the reflectivity of the target may strongly depend on the polarization orientation of the illuminating waves, and somehow, the target structure acts as a more or less polarization-sensitive filter. This may require adaptation of polarization to maximize target reflectivity. The polarization orientation of the illuminating wave can be controlled by proper weighting and simultaneous transmission of the waveform from two polarimetric Tx antenna ports. With phase control included, circular or elliptic polarization can be generated from two orthogonal linear polarized antenna ports. The same adaptation principle also applies to the Rx chain. Note that usual definitions of linear polarization as Vertical (V) and Horizontal (H) apply to the local coordinate system of Tx and Rx antennas, which may not be aligned in H and V between each other. The same applies to the local coordinate system of the target, and also does not apply to the Earth-centered coordinate system. This means that just simplifying to vertical (VV) or horizontal (HH) polarization at the transmitter (Tx) and receiver (Rx) is not enough in most cases if the Tx and Rx antennas are randomly oriented in space, as it may be the case if agile ISAC sensor platforms are included.
@@ V-F Spatial Precoding, Beamforming, and Bidirectional Estimation
Spatial precoding is a well-established method in mobile radio that came into play with the introduction of multiple antennas. It helps increase link performance, reduce interference, host and address multiple users, and increase data rate and bandwidth efficiency. Spatial precoding will also be very important for ISAC, especially for MS-ISAC [BF1]. Since 5G NR uses a flexible architecture based on codebook and non-codebook (full adaptive) schemes, spatial radio access can be controlled for omnidirectional or directive sensing. Without going too much into the technical details, we should emphasize that spatial precoding, especially beamforming, becomes more important at higher frequencies (FR2), where the need for higher beamforming gain increases because of link budget issues. However, hybrid (mixed analog and digital) antenna interfaces seem to be the dominating solution, where digitization takes place at the output of the beamformer, not at the antenna ports. Multiple simultaneous beams for the same user are also possible. 5G NR allows transmit precoding and receive combining, and also supports beamforming in the UL, which means that the gNB can assign a spatial precoding matrix to the UE, which can be a dedicated device acting as a sniffer sensor, as described above. In the sequel, we will discuss some striking applications of beamforming and spatial access for MS-ISAC, though we do not claim to be exhaustive.
@@
Beamforming. The initial purpose of beamforming is to enhance the SNR (and consequently the range), as it is associated with an antenna gain. In the context of distributed cooperative MS-ISAC networks, one application is to balance the SNR to achieve smooth coverage and performance distribution irrespective of the density of the spatial distribution of the sensor nodes. MS-ISAC network densification will also go along with a proper choice of the frequency. Higher frequencies will be more appropriate for sensing hotspots since the wider bandwidth at FR2 offers more resolution capabilities in delay. But, because of the Friis and radar equations, higher frequencies are more subjected for transmission loss since the effective antenna area is becoming smaller for the same gain. This can be well compensated for by antenna arrays. However, this is accompanied by directivity. Hence, we would need time-consuming angular scanning in the surveillance cycle. Moreover, it will be appropriate to allocate different beams for the target and for the UE direction. Multifrequency sensing can help to get a fast (quick and dirty) overview picture of the scene and some first indication of possible targets using wide-open beam antennas at lower frequencies that do not require expensive angular scanning. The result is then used to instruct the more selective FR2 beamformer in a second step [ref83].
@@
Beamforming directivity can be applied deliberately to reduce interference and multipath clutter through spatial filtering, especially when executed in a bidirectional manner at both the Tx and Rx. The generic antenna architecture depicted in Figure 22 is circular on both sides, making it ideal for symmetric D2D scenarios, such as V2V, with arbitrary spatial orientation of the devices. An equivalent antenna setup can be found for V2I communication, where planar arrays at the infrastructure side may be more appropriate. The direct LOS link is important for data transfer and for obtaining the correlation reference. LOS beamforming suppresses multipath, which supports transmit signal recovery. A second surveillance channel points to the target, while clutter 1 is not detected by both beams and is therefore not visible. The case of target clutter 2 is particularly interesting. Reflections from this object could not be filtered out by the Tx beam, but can be well suppressed by the Rx beam pointing towards Tx. Therefore, the Rx-LOS beam is important for receiving a clean, multipath-free copy of the transmitted signal. When the same object is considered as the target, the corresponding Rx beam pointing to the target makes it visible and also reduces LOS interference. This example shows how the two bidirectional beams support each other and may increase directive filtering a lot. However, it also increases beam search effort. It is worth mentioning that, in a 3D geometry, matching the two opposite beams must also include polarization orientation, especially when azimuth and elevation matter and the orientation of the antennas may be more or less arbitrary [ref68]. Since the two beams are served by the same radio interface, mutual coherence is maintained, which is important for eToF and eDoppler estimation in CPCL.
@@
Figure 22: Bidirectional beamforming supporting clean LOS filtering and clutter suppression.
High resolution DoD/DoA estimation. Beamforming can be very important to support ISAC radio access and also provide some directional information about impinging waves, especially if the beam is narrow. However, it does not directly allow high-resolution directional estimation beyond the Rayleigh limit. The decisive factor for high-resolution DoA estimation is phase difference at displaced antenna outputs. This would enable a much more accurate DoA estimation than simply selecting the radiation pattern or relying on magnitude differences for DoA interpolation within the main beam lobe. The problem is that the analog / hybrid 5G beamforming interface hides the antennae in / outputs for individual access and processing, which prevents the estimation phase differences. However, high-resolution estimation schemes such as MUSIC or Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT) would require mutual antenna output correlation, which is a quadratic (hence nonlinear) operation and not linear beamforming for building the array correlation matrix.
@@
In the case of hybrid beamforming, the outputs of the analog multibeam network map a huge number of hundreds or even thousands of physical antennas by a linear weight-and-sum network to a small number of virtual antenna ports where digitization takes place. [ref97], for instance, describes the beamformer design based on limited-resolution phase shifters. In the following, we introduce an approach for high-resolution DoA estimation that makes intelligent use of the precoding beamforming vectors. We assume that the beam patterns overlap and the beam outputs explicitly differ in phase, which can then be used for DoA estimation. The required number of virtual antenna ports is relatively small (much smaller than the number of antennas). This is because we only need to resolve those paths by angle that are not resolved by other parameters (delay, Doppler, or the transmit beamformer).
@@
A simple example is given in Figure 23. There are two beams formed, both of which have the same gain pattern. The beamforming vector is effectively shifted by one antenna element, which emulates two identical, overlapping subarrays effectively shifted by one antenna element. This results in so-called translation-invariant radiation patterns, meaning that any impinging wave is weighted by the same gain pattern. However, as the two beams have different phase centers, they are very well suited for high-resolution DoA estimation. The well-known monopulse technique for high-resolution DoA estimation in radar is reminiscent of a simple replica of the same principle in the analog world. On a more recent time scale, the two arrays directly correspond to the shift-invariant pair of arrays that forms the basic principle of the ESPRIT algorithm. Obviously, the proposed approach can be further modified by increasing the number of subarrays and varying the phase center distances. Multiple overlapping shift-invariant subarrays would allow resolving multiple wave directions by subarray smoothing and not only increasing the accuracy of DoA estimation for one impinging path. This method can be considered as a hybrid implementation of the ESPRIT algorithm. It is limited to linear and planar arrays, which is not a serious drawback since integrated arrays primarily come as planar structures. Another advantage is that it does not require a specific array design with very sharp pencil beams. Since DoA estimation accuracy is gained by phase comparison, it is more important that the phase shifters are identical. Resolution of multiple paths within the same common beam is possible by exploiting the phase differences. On the other hand, beamforming can suppress waves from outside of the uniquely resolvable sector. This is indicated in Figure 23 by the narrower beam. This would allow the distance between antenna elements to be extended beyond the half-lambda limit, which would increase estimation accuracy.
@@
Even if the goal is to estimate the DoD/DoA based on the maximum likelihood parameter estimation framework [ref64, ref67], the radiation pattern design can be relaxed since the specific pattern is not as important as long as the pattern is known. Therefore, we need to know how the complex pattern transforms from the specified coefficients to the radiated field to enable the iterative cancellation of overlapping beam patterns. Thus, the focus shifts from optimizing quantized coefficients as in [ref97] to calibrating the radiation pattern. Lastly, estimation of DoD requires identifying the waveforms transmitted via the different antenna ports at the receive (Rx) side. Therefore, some kind of orthogonalization in the frequency-time frame is necessary.
@@
Figure 23: Emulating two shift-invariant radiation patterns for high resolution direction finding. The wider beams allow covering a wide angular sector, while the narrower beams reduce the angular field of view by spatial filtering.
@@
Multidimensional estimation. The multidimensionality aspect of parameter estimation refers to the bistatic propagation scenario, which includes the bidirectional geometry (in azimuth and elevation at Tx and Rx), ToF and Doppler for each propagation path, weighted with a $2\times 2$ complex-valued polarimetric Jones matrix. Thus, the bidirectional estimation extends the ToF-Doppler estimation described in the subsection on Model-based Bistatic Target Parameter Estimation to a joint 6D estimation procedure. When performing DoADoD estimations, impairments of antennas and antenna arrays can be of highest importance as they are most susceptible to RF effects. In this case, the device data model will be determined by calibration measurements of the installed radiation patterns. The traditional approach of model-based estimation decomposes into a local gradient-based and a global parameter optimization procedure with constraints. Actual research challenges address efficient real-time implementation in networks with distributed computational resources and AI support to multi-domain convex optimization with detection of relevant propagation paths. A related issue is in situ training, which effectively updates calibration throughout operation [ref67, schieler2025prop_est_cnn]. Finally, advanced AI-based methods can be used to adjust the model parameters, as this can be a challenging problem of non-convex optimization, which can become even more complex when device data from calibration is not perfect or not fully available from calibration measurements in reference setups [ref35], but has to be learned implicitly by ML in operational situations.